# Yuyang Tao, Shufei Ge

### v1

@@ 2.4 Evaluation metric
The silhouette coefficient ($SC$) is often used as a measure to evaluate the quality of (overlap) clustering hanClusterAnalysis2012 . The $SC$ assesses how well the clusters are separated, how close the clusters are, and is stable for overlap clustering. Suppose we have $n$ samples in dataset $D$ that could be divided into $k$ clusters: $C_{1},...,C_{k}$. For a data point $x\in D$, the compactness of the cluster to which $x$ belongs can be defined as:
$a(x)=\frac{\sum_{x^{\prime}\in C_{i},x\neq x^{\prime}}d(x,x^{\prime})}{|C_{i}|-1},$
@@
$b(x)=\mathop{\min}_{C_{j}:j\neq i}\bigg{\{}\frac{\sum_{x^{\prime}\in C_{j}}d(x,x^{\prime})}{|C_{j}|}\bigg{\}},$
@@
The value of the $SC$ ranges from $-1$ to $1$. A value close to 1 indicates the point is close to the current cluster, and more distinct from other clusters. If the $SC$ is less than 0, it means the point is closer to other clusters compared to the current cluster, and it usually indicates bad clustering results. The clustering of the whole dataset can be assessed by taking average of the $SC$s of all points.
@@
The $SC$ reflects the clustering quality of data points, but it does not evaluate the topology structure of a Mapper output graph. Figure 2 shows an example of a Mapper graph that has high $SC$ but poor topological structure. The left panel has higher $SC$ but poor topological feature while the right panel has lower $SC$ but good topological structure. Therefore, a metric that evaluates both clustering and topological structure is needed. Although, the topological information encoded in Mapper is well studied theoretically deyTopologicalAnalysisNerves2017 . There is no practical metric to evaluate the topology structure of Mapper graph. Most existing work evaluates the topology structure of Mapper graphs based on the clustering results only. The extended persistence diagram has been proven to be a powerful tool for capturing topological signatures of Mapper graph carriereStructureStabilityOneDimensional2018 . In this manuscripts, we alleviate it to provide a simple quantitative metric for evaluating the topology structure of Mapper output.
@@
Figure 2: An example of the classical Mapper algorithm has higher $SC$ but poor topological structure, the dataset is shown in Figure 5 (a), the classical Mapper is implemented on this dataset. (a) A graph produced by the classic Mapper with parameters: $n=12,p=0.01$. $SC:0.283$, $SC_{adj}:0.521$. (b) A graph produced by the classic Mapper with parameters: $n=12,p=0.1$, $SC:0.246$, $SC_{adj}:0.812$.
We shall introduce some background of the extended persistence diagram, which are detailed in carriereStructureStabilityOneDimensional2018 . The extended persistence diagram is an extension of the persistence theory. The extended persistence theory uses both sublevel sets and superlevel sets as a filtration to provide more information than the original persistence theory. Like the persistence diagram, the extended persistence diagram spans by points of birth time and death time. As a result, points on the extended persistence diagram can be located anywhere on the plane, unlike the ordinary persistence diagram where points can only be strictly above the diagonal.
Definition 5 (Extended persistence diagram).
The extended persistence diagram is a multiset of points in the Euclidean plane $\mathbb{R}^{2}$. For a graph $G=(V,E)$ and a function defined on its nodes $f:V\rightarrow\mathbb{R}$, the extended persistence diagram is computed by both sublevel sets and superlevel sets, denoted as $Dg(G,f)$. According to the combination of sublevel sets and superlevel sets filtration, each point in the diagram is classified as either $Ord_{0}$, $Rel_{1}$, $Ext^{+}_{0}$ or $Ext^{-}_{1}$.
@@
Briefly speaking, given a graph and a function defined on its nodes, we can compute the extended persistence diagram that reflects the topological features of this graph. For a Mapper graph, the function on a node can be naturally defined as the mean value of points in the node. The points on the diagram can be regarded as signatures of the Mapper graph. However, these signatures are not always meaningful, points near the diagonal are noises. To distinguish noise and real signal, one simple but effective way is to compute the confidence set by bottleneck bootstrap and separate noise and signal based on this confidence set carriereStatisticalAnalysisParameter2017 . Bottleneck bootstrap is an effective way to compute confidence set on a persistence diagram fasyConfidenceSetsPersistence2014 . It uses bootstrap samples to get a bootstrap Mapper graph and then calculate the bottleneck distance from the original Mapper graph. Repeat this step many times, we can obtain an approximate distribution of bottleneck distance. With this distribution we can easily calculate the confidence set.
@@
We introduce a metric to evaluate the quality of the extended persistence diagram dependent on the confidence set, we call this metric the topological signal rate (TSR). The TSR is defined as the number of real signal points divided by the total number of points on the extended persistence diagram, serving as a quantitative indicator of the quality of the extend persistence diagram.
@@
The topological signal rate is a scalar, it evaluates the quality of a persistence diagram or extended persistence diagram, denoted as TSR:
@@
where $N$ is the number of all points on the diagram and $N_{signal}$ is the number of topological signal on the diagram.
An example of an extended persistence diagram and its TSR is illustrated in Figure 3. The gray area is the confidence set estimated by the bottleneck bootstrap. Points inside are noises and those outside are signal points. The TSR in this example is 0.25, indicating poor quality of the corresponding Mapper graph. To get a proper evaluation of the Mapper graph, we can combine the TSR with $SC$ through weighted averaging. We call this metric as adjusted silhouette coefficient.
Figure 3: An example of the extended persistence diagram. There are totally 8 points in the diagram, the gray area is computed by the bottleneck bootstrap, points inside this area are noise. Thus, 6 points are noises and 2 points are signals, the TSR is 0.25 .
@@
The adjusted silhouette coefficient is the weighted average of the normalized $SC$ and the $TSR$, denoted as $SC_{adj}$, this value can serve as a metric to evaluate Mapper type algorithms:
@@
Figure 2 compares the value of $SC_{adj}$ with the original $SC$. The original $SC$ overlooks the topological structure of Mapper graph, while the $SC_{adj}$ can give a reasonable evaluation. In the next section, we show more examples of $SC_{adj}$ and validate its effectiveness through some comparisons.
3 Simulation
In this section, we compare our proposed D-Mapper algorithm with the classic Mapper algorithm via several experiments. In this paper, we implement D-Mapper by expanding the Mapper algorithm in the KeplerMapper package version 2.0.1 KeplerMapper_JOSS in python version 3.11.0. The extended persistence diagram and bottleneck bootstrap are computed by a python package GUDHI version 3.8.0 gudhi:urm .
We compare our proposed D-Mapper algorithm with the classic Mapper algorithm using the metric $SC_{adj}$ and $SC_{norm}$. We set the interval number $n$ to be identical for both algorithms and use the same clustering algorithm within each hypercube. The clustering method is the density-based spatial clustering of applications with noise (DBSCAN) implemented in the scikit-learn library version 1.1.3. We use a fixed grid to find the best model concerning the $SC_{adj}$ for both the D-Mapper and the classic Mapper. $50$ equally spaced grids in range $(0,\alpha_{max})$ are used to select parameter $\alpha$ in the D-Mapper. Similarly, $50$ equally spaced grids in range $(0,0.5)$ are used to tune the overlap percentage $p$ in classic Mapper algorithm. The bottleneck bootstrap sampling steps are set to $100$ to compute $85$ percent confidence regions. In the next section, we compare the D-Mapper with the classic Mapper on several examples.
@@ 3.1 Two disjoint circles
In this experiment, we choose the filter function to be a function that projects the original data onto the X-axis. We set the number of intervals $n$ to $12$. The parameters $p$ of the classic Mapper algorithm and parameter $\alpha$ of the D-Mapper are tuned via grid search, $n=12,p=0.02$, $n=12,\alpha=0.127$. The comparison of different evaluation metrics is shown in Table 1 and Figure 4. In Figure 4, the color of a node indicates the average value of projected data in the node. The output graphs of the classic Mapper and D-Mapper are similar, and both algorithms capture the topological features of the dataset effectively without noise, result in a TSR of 1 for both algorithms. The D-Mapper performs better than the classic Mapper in terms of clustering, as evidenced by its higher $SC_{norm}$. This indicates that D-Mapper has an advantage in identifying more meaningful clusters. Moreover, we show cases that both Mapper algorithms concerned on $SC$ only in Figure 4 (c) and (d). In these two cases the $SC_{norm}$ are higher but the TSR are lower than the results concerned on $SC_{adj}$. The topology structures of these two cases are obviously different from the dataset in these two figures. The distinct results in this example validate the utility of our proposed metric. It provides a quantitative approach for evaluating Mapper-type algorithms from both topological signal preservation and clustering.
Table 1: Results of the D-Mapper and classic Mapper on two disjoint circles dataset.The $1$st and $3$rd rows are results that with the larger $SC_{adj}$. The $2$nd and $4$th rows are cases when the output graphs have larger $SC$ values, but lower TSRs.
[…]
@@
Figure 4: Results of the classic Mapper and D-Mapper on the two disjoint circles. (a) The output graph of the classic Mapper with the largest $SC_{adj}$ (the $1$st row of Table 1): $n=12,p=0.02$. (b) The output graph of D-Mapper with the largest $SC_{adj}$ (the $3$rd row of Table 1): $n=12,\alpha=0.127$. (c) An example produced by the D-Mapper with larger $SC$ but lower TSR (the $2$nd row of Table 1): $n=12,\alpha=0.159$. (d) An example produced by the classic Mapper with larger $SC$ but lower TSR (the $4$th row of Table 1): $n=12,p=0.005$.
@@ 3.2 Two intersecting circles
In this section, we compare the performance of the D-Mapper and the classic Mapper on a two intersecting circles dataset shown in Figure 5 (b). The data points are generated from two intersecting circles with a radius of 1 and centers (0,0) and (1.5,0), respectively. The data generating process is similar to the previous example, 5000 points are sampled from each circle. The filter function is the same as in the previous example. For the classic Mapper algorithm, the number of intervals $n$ is set to $8$ and the overlap rate $p$ is chosen as $0.02$. As for the D-Mapper algorithm, the number of intervals is the same as the classic Mapper’s and the parameter $\alpha$ is tuned to $0.088$. The result of this experiment is given in Figure 6 and Table 2. The D-Mapper outperforms the classic Mapper concerned on $SC$ scores, and has the same topological signal rate as the classic Mapper. As a result, the $SC_{adj}$ of D-Mapper is higher than the classic Mapper. Similar with the two disjoint circles example, Figure 6 (c) and (d) also show cases that when the metric $SC$ fails.
Table 2: Results of the D-Mapper and classic Mapper on two disjoint circles dataset. The $1$st and $3$rd rows are results that with the larger $SC_{adj}$. The $2$nd and $4$th rows are cases when the output graphs have larger $SC$ values, but lower TSRs.
[…]
@@
Figure 6: Results of the classic Mapper and D-Mapper on the two intersecting circles. (a) The output graph of the classic Mapper with the largest $SC_{adj}$ (the $1$st row of Table 2): $n=8,p=0.02$. (b) The output graph of D-Mapper with with the largest $SC_{adj}$ (the $3$rd row of Table 2): $n=8,\alpha=0.088$. (c) An example produced by the D-Mapper with larger $SC$ but lower TSR (the $2$nd row of Table 2): $n=8,\alpha=0.12$. (d) An example produced by the classic Mapper with larger $SC$ but lower TSR (the $2$nd row of Table 2): $n=8,p=0.02$.
Both the results of the two disjoint circles and the two intersecting circles indicate that our proposed metric $SC_{adj}$ is more stable than the metric $SC$ measuring both the quality of overlap clustering and extended persistence homology of output of Mapper type algorithm.
@@
In this experiment, we compare the D-Mapper algorithm with the classic Mapper algorithm on a 3D cat dataset with a more complex topology structure. The 3D cat dataset is originally created by 2004Deformation . Figure 7 (a) shows a visualization of the 3D cat dataset. The filter function returns the averaged pairwise distance between the points and all other points. The parameters of the classic Mapper are set to $n=9,p=0.32$; and the parameters of the D-Mapper are set to $n=9,\alpha=0.01$. The clustering method is DBSCAN with a radius of $0.1$ and a minimum of samples $5$. The results and evaluation metrics are shown in Figure 7 and Table 3. The D-Mapper algorithm results higher $SC_{norm}$ scores and the same TSR compares to the classic Mapper algorithm. The output graph of the D-Mapper has one more loop than the classic Mapper, which represents the ears of cats (yellow dots in Figure 7 c-d). This example also shows that the D-Mapper could capture subtle features of complex objects better than the classic Mapper.
@@
Figure 7: Results of the classic Mapper and D-Mapper on the 3D cat dataset. (a) Visualization of the 3D cat dataset. (b) The intervals produced by both Mappers. The intervals of both Mappers are very close except intervals near to 1, this difference gives more details on cat head in D-Mapper. (c) The graph of the classic Mapper with the largest $SC_{adj}(0.740)$: $n=9,n=0.32$. (d) The graph of D-Mapper with the largest $SC_{adj}(0.755)$: $n=9,\alpha=0.01$.
3.4 Summary of experiments
In all experiments, the D-Mapper outperforms the classic Mapper algorithm concerning on the metric $SC_{adj}$. The $SC_{norm}$ of the D-Mapper are all higher than the classic Mapper algorithm, and all TSRs are $1$. This indicates that the D-Mapper can achieve better clustering than the classic Mapper algorithm while outputing high quality reeb graph approximations. Note that we tune parameters by grid search the highest $SC_{adj}$, thus the TSR is relatively high. In many cases the TSR could be low, as we show in Table 1 and Table 2. Our experimental results also indicate that our proposed metric $SC_{adj}$ is more stable than the metric $SC$.
4 Application

### v2

@@ 2.4 Evaluation metric
The silhouette coefficient ($SC$) is often used as a measure to evaluate the quality of (overlap) clustering [17]. The $SC$ assesses how well the clusters are separated, how close the clusters are, and is stable for overlap clustering. Suppose we have $n$ samples in dataset $D$ that could be divided into $k$ clusters: $C_{1},...,C_{k}$. For any two data points $x,~{}x^{\prime}\in D$, the compactness of the cluster to which $x$ belongs can be defined as:
$a(x)=\frac{\sum_{x^{\prime}\in C_{i},x\neq x^{\prime}}d(x,x^{\prime})}{|C_{i}|-1},$ (1)
@@
$b(x)=\mathop{\min}_{C_{j}:j\neq i}\bigg{\{}\frac{\sum_{x^{\prime}\in C_{j}}d(x,x^{\prime})}{|C_{j}|}\bigg{\}},$ (2)
@@
The value of the $SC$ ranges from $-1$ to $1$. A value close to $1$ indicates the point is close to the current cluster, and more distinct from other clusters. If the $SC$ is less than 00, it means the point is closer to other clusters compared to the current cluster, and it usually indicates bad clustering results. The clustering of the whole dataset can be assessed by taking the average of the $SC$s of all points.
@@
The $SC$ reflects the clustering quality of data points, but it does not evaluate the topology structure of a Mapper output graph. Figure 2 shows an example of different Mapper graphs on the same dataset. The left panel has a higher $SC$ but a poor topological feature. Conversely, the right panel has a lower $SC$ but a good topological structure. Figure 2 (a) shows that small overlap $p$ makes one of the circles disconnected, however, more separable clusters lead to a high $SC$. Therefore, a metric that evaluates both clustering and topological structure is needed. Although, the topological information encoded in Mapper has been well studied theoretically [15]. There is no practical metric to evaluate the topology structure of Mapper graphs. Most existing work evaluates the topology structure of Mapper graphs based on the clustering results only. The extended persistence diagram has been proven to be a powerful tool for capturing topological signatures of Mapper graph [18]. In this manuscript, we alleviate it to provide a simple quantitative metric for evaluating the topology structure of Mapper output.
@@
Figure 2: An example of the classic Mapper algorithm on the same dataset but outputs different graphs. The dataset is shown in Figure 6 (a), the classical Mapper is implemented, and the clustering algorithm is DBSCAN with a radius of $0.5$ and a minimum of samples $3$. (a) The output graph when $n=12,p=0.01$ and its $SC=0.283$, $SC_{adj}=0.521$. This graph has a higher $SC$ but a poor topological structure. (b) The output graph when $n=12,p=0.1$ and its $SC=0.246$, $SC_{adj}=0.812$. This graph has a lower $SC$ but a good topological structure.
@@
The extended persistence diagram is an extension of the persistence diagram, the extended persistence diagram uses both sub-level sets and super-level sets as a filtration, therefore it contains more information than the original persistence diagram [19]. In practical, given a graph with a function defined on its nodes, we can compute the extended persistence diagram that reflects the topological features of this graph. For a Mapper graph, the function on a node can be naturally defined as the mean value of points in the node. The points on the diagram can be regarded as signatures of the Mapper graph. However, these signatures are not always meaningful, points near the diagonal are often seen as noise. To distinguish the real signal from the noise, one simple yet effective method is to calculate the confidence interval using bottleneck bootstrap and then identify the noise based on this interval[9]. Bottleneck bootstrap is an effective way to compute confidence set on a persistence diagram [20]. It uses bootstrap samples to get a bootstrap Mapper graph and then calculates the bottleneck distance from the original Mapper graph. Repeat this step many times, and an approximate distribution of bottleneck distances can be obtained, then we can easily calculate the confidence set with this approximate distribution. More details of the bottleneck bootstrap algorithm are shown in Appendix B.
@@
We introduce a metric to evaluate the quality of the extended persistence diagram depending on the confidence set, we call this metric the topological signal rate, denoted as $TSR$. The $TSR$ is defined as the number of real signal points divided by the total number of points on the extended persistence diagram, serving as a quantitative indicator of the quality of the extend persistence diagram.
@@
The topological signal rate is a scalar, it evaluates the quality of a persistence diagram or extended persistence diagram, denoted as $TSR$,
@@
where $N$ is the number of all points on the diagram and $N_{signal}$ is the number of topological signals on the diagram.
An example of an extended persistence diagram and its $TSR$ is illustrated in Figure 3. The gray area is the confidence set estimated by the bottleneck bootstrap. Points inside are noises and those outside are signal points. The $TSR$ in this example is $0.25$, indicating poor quality of the corresponding Mapper graph. To get a proper evaluation of the Mapper graph, we can combine the $TSR$ with $SC$ through weighted averaging. We call this metric an adjusted silhouette coefficient.
Figure 3: An example of the extended persistence diagram. There are a total of $8$ points in the diagram, the gray area is computed by the bottleneck bootstrap, and points inside this area are noise. Thus, $6$ points are noises and $2$ points are signals, the $TSR$ is $0.25$.
@@
The adjusted silhouette coefficient is the weighted average of the normalized $SC$ and the $TSR$, denoted as $SC_{adj}$, this value can serve as a metric to evaluate the Mapper-type algorithms:
@@
Figure 2 compares the value of $SC_{adj}$ with the original $SC$. The original $SC$ overlooks the topological structure of Mapper graph, while the $SC_{adj}$ can give a reasonable evaluation. In the next section, we show more examples to validate the effectiveness of $SC_{adj}$.
3 Results
In this section, we compare our proposed D-Mapper algorithm with the classic Mapper algorithm via several experiments. In this paper, we implement D-Mapper by expanding the Mapper algorithm in the KeplerMapper package version 2.0.1 [21] in Python version 3.11.0. The extended persistence diagram and bottleneck bootstrap are computed by a Python package GUDHI version 3.8.0 [22].
We compare our proposed D-Mapper algorithm with the classic Mapper algorithm using the metric $SC_{adj}$ and $SC_{norm}$. We set the interval number $n$ to be identical for both algorithms and use the same clustering algorithm within each hypercube. The clustering method is the density-based spatial clustering of applications with noise (DBSCAN) implemented in the scikit-learn library version 1.1.3. We use a fixed grid to find the best model concerning the $SC_{adj}$ for both the D-Mapper and the classic Mapper. $50$ equally spaced grids in range $(0,\alpha^{\prime})$ are used to select parameter $\alpha$ in the D-Mapper. Similarly, $50$ equally spaced grids in range $(0,0.5)$ are used to tune the overlapping percentage $p$ in the classic Mapper algorithm. The bottleneck bootstrap sampling steps are set to $100$ to compute $85$ confidence level. In this section, we compare the D-Mapper with the classic Mapper via several examples.
@@ 3.1 Two disjoint circles
In this experiment, we choose the filter function to be a function that projects the original data onto the $X$-axis. We set the number of intervals $n$ to $12$. The clustering algorithm is DBSCAN with a radius of $0.5$ and a minimum of samples $3$. The parameters $p$ of the classic Mapper algorithm and parameter $\alpha$ of the D-Mapper are tuned via grid search, $p=0.02$ and $\alpha=0.127$. The comparison of different evaluation metrics is shown in Table 1 and Figure 4. In Figure 4, the color of a node indicates the average value of the projected data within the node. The output graphs of the classic Mapper and D-Mapper are similar, and both algorithms capture the topological features of the dataset effectively without noise, result in a $TSR$ of 1 for both algorithms. The D-Mapper performs better than the classic Mapper in terms of clustering, as evidenced by its higher $SC_{norm}$. This indicates that D-Mapper has an advantage in identifying more meaningful clusters. Moreover, we show cases that both Mapper algorithms concerned on $SC$ only in Figure 4 (c) and (d). In these two cases the $SC_{norm}$ are higher but the $TSR$ are lower than the results concerned on $SC_{adj}$. The topology structures of these two cases are obviously different from the dataset in these two figures. The distinct results in this example validate the utility of our proposed metric. It provides a quantitative approach for evaluating Mapper-type algorithms from both topological signal preservation and clustering.
Table 1: Results of the D-Mapper and classic Mapper on two disjoint circles dataset.The $1$st and $2$nd rows are results that with the larger $SC_{adj}$. The $3$rd and $4$th rows are cases when the output graphs have larger $SC$ values, but lower $TSR$s.
Algorithm $SC_{norm}$ $TSR$ $SC_{adj}$
Classic Mapper 0.640 1.00 0.820
D-Mapper 0.716 1.00 0.858
@@
Figure 4: Results of the classic Mapper and D-Mapper on the two disjoint circles. (a) The output graph of the classic Mapper with the largest $SC_{adj}$ (the $1$st row of Table 1): $n=12,p=0.02$. (b) The output graph of the D-Mapper with the largest $SC_{adj}$ (the $2$nd row of Table 1): $n=12,\alpha=0.127$. (c) An example produced by the classic Mapper with larger $SC$ but lower $TSR$ (the $3$rd row of Table 1): $n=12,p=0.005$. (d) An example produced by the D-Mapper with larger $SC$ but lower $TSR$ (the $4$th row of Table 1): $n=12,\alpha=0.159$.
@@ 3.2 Two intersecting circles
In this section, we compare the performance of the D-Mapper and the classic Mapper on a two intersecting circles dataset shown in Figure 6 (b). The data points are generated from two intersecting circles with a radius of $1$ and centers $(0,0)$ and $(1.5,0)$, respectively. The data generating process is similar to the previous example, $5000$ points are sampled from each circle. The filter function and clustering algorithm is the same as in the previous example. The number of intervals $n$ is set to $8$. The parameters $p$ of the classic Mapper algorithm and parameter $\alpha$ of the D-Mapper are tuned to $p=0.02$ and $\alpha=0.088$. The results are given in Figure 5 and Table 2. As shown in Figure 5 (a,b), the D-Mapper outperforms the classic Mapper concerned on $SC$ scores (or the $SC_{adj}$) and has the same topological signal rate as the classic Mapper. Similar to the two disjoint circles example, Figure 5 (c) and (d) also give cases when the metric $SC$ fails to work.
Table 2: Results of the D-Mapper and classic Mapper on two intersecting circles dataset. The $1$st and $2$nd rows are results that with the larger $SC_{adj}$. The $3$rd and $4$th rows are cases when the output graphs have larger $SC$ values, but lower $TSR$s.
Algorithm $SC_{norm}$ $TSR$ $SC_{adj}$
Classic Mapper 0.574 1.00 0.787
D-Mapper 0.640 1.00 0.820
@@
Figure 5: Results of the classic Mapper and D-Mapper on the two intersecting circles. (a) The output graph of the classic Mapper with the largest $SC_{adj}$ (the $1$st row of Table 2): $n=8,p=0.02$. (b) The output graph of the D-Mapper with the largest $SC_{adj}$ (the $2$nd row of Table 2): $n=8,\alpha=0.088$. (c) An example produced by the classic Mapper with larger $SC$ but lower $TSR$ (the $3$rd row of Table 2): $n=8,p=0.02$. (d) An example produced by the D-Mapper with larger $SC$ but lower $TSR$ (the $4$th row of Table 2): $n=8,\alpha=0.12$.
Both the results of the two disjoint circles and the two intersecting circles indicate that our proposed metric $SC_{adj}$ is more stable than the metric $SC$ measuring both the quality of overlap clustering and extended persistent homology of output of Mapper type algorithm.
@@
In this experiment, we compare the D-Mapper algorithm with the classic Mapper algorithm on a 3D cat dataset with a more complex topology structure. The 3D cat dataset is originally created by [23]. Figure 7 (a) shows a visualization of the 3D cat dataset. The filter function is the sum of all coordinates. The parameters of the classic Mapper are set to $n=9,~{}p=0.32$; and the parameters of the D-Mapper are set to $n=9,~{}\alpha=0.01$. The clustering method is DBSCAN with a radius of $0.1$ and a minimum of samples $5$. The results and evaluation metrics are shown in Figure 7 and Table 3.
The D-Mapper algorithm results higher $SC_{norm}$ scores and the same $TSR$ compares to the classic Mapper algorithm. The output graph of the D-Mapper has one more loop than the classic Mapper, which represents the relationship between the ears, face (including jaw) and neck of the cat (yellow, olive green and light green dots in Figure 7 c-d). This example also shows that the D-Mapper could capture subtle features of complex objects better than the classic Mapper due to the flexible intervals and overlap ratios.
@@
Figure 7: Results of the classic Mapper and D-Mapper on the 3D cat dataset. (a) Visualization of the 3D cat dataset. (b) The intervals produced by both Mappers. The intervals of both Mappers are very close except intervals near to $1$, this difference gives more details on cat head in D-Mapper. (c) The graph of the classic Mapper with the largest $SC_{adj}(0.740)$: $n=9,p=0.32$. (d) The graph of the D-Mapper with the largest $SC_{adj}(0.755)$: $n=9,\alpha=0.01$.
3.4 Covid-19 dataset